### v2

@@ Methods
Next, to adapt the classifier to also predict class labels for “noisy” data, we froze all single-site terms involving non-mask tokens as well as any pairwise interaction terms between two non-mask tokens (i.e. we only further trained single site parameters corresponding to mask tokens and trained pairwise interaction terms involving at least one mask token). For each gradient step, we first sampled the entire un-noised dataset, then for each sequence in the dataset we randomly sampled a time between 00 and $1$, and masked the sequence by sampling from the forward noising process up to that time. The model was again trained for $400$ epochs without any regularization, at which point the loss function converged. Training the model in this way, wherein the parameters for the non-mask terms are frozen, has the nice property that the noisy model is gauranted to be equivalent to the non-noisy model at $t=1$.
Second-round guided ABE library design
We generated the second-round library by guiding ESM3 and FMIF respectively with the classifier we trained based on the experimental data collected on sequences from the first round. We used the same sampling hyperparameters for each model as we did in the first round, and a guidance factor $\gamma=100$. We designed 944 sequences with each model. All of the generated sequences had a predicted probability of being highly active greater than 0.5. We also included 112 sequences from the first-round library as controls so we could calibrate the log-enrichment measurements between the two rounds. We selected these 112 control sequences with by sampling uniformly across the range of observed log enrichment in the first round.
The 1,888 designed sequences and 112 control sequences were assayed and sequenced with the same protocols as were done in the first round (Figure 9). To enable direct comparison between the first- and second-round libraries, we included 112 sequences from the first round as controls in the second-round library (Figure 10a). These control sequences were selected to span the full range of observed log enrichment values from the first round, providing calibration points across the entire activity spectrum. This design allowed us to quantify and correct for batch effects between the two experimental rounds. We used these control sequences to learn a linear transformation that maps log enrichment values from the second round to their corresponding values in the first round. Specifically, we fit a linear regression model:
$\log e_{\text{round 1}}=\alpha\cdot\log e_{\text{round 2}}+\beta$ (33)
@@
To assess the uncertainty in this calibration procedure, we employed a bootstrap approach using the 112 control sequences. Specifically, we performed 1,000 bootstrap iterations, each time randomly sampling 112 sequences with replacement from our control set and fitting a linear regression model. For each bootstrap sample, we derived a unique set of mapping parameters ($\alpha_{i}$, $\beta_{i}$) and used these to transform the entire second-round dataset. We then calculated the median log enrichment value of the transformed second-round sequences for each bootstrap iteration, generating a distribution of potential outcomes that reflects the inherent variability in our experimental measurements. From this distribution of bootstrap outcomes, we identified the 5th and 95th percentile mappings based on the median log enrichment values. These percentiles represent conservative and optimistic estimates, respectively, of the transformation from round 2 to round 1 log enrichment values (Figure 10b).
Appendix A
Supplementary tables
Table 1: Structure-conditioned protein sequences generation with stability guidance results for all proteins evaluated by predicted stability and folding into correct structure. We defined success as sequences that were generated with RMSD $\leq 2$Å and $\Delta\Delta G\leq 0$. $p_{\theta}(\mathbf{s}\,|\,\mathbf{x})$ Temp is the temperature of the baseline structure-conditioned model without stability guidance, while parenthetical $\gamma$ denotes the guidance strength.

### v4

@@ 1.11.6 Predicting ABE activity from the first-round library
Next, to adapt the classifier to also predict class labels for “noisy” data, we froze all single-site terms involving non-mask tokens as well as any pairwise interaction terms between two non-mask tokens (i.e. we only further trained single site parameters corresponding to mask tokens and trained pairwise interaction terms involving at least one mask token). For each gradient step, we first sampled the entire un-noised dataset, then for each sequence in the dataset we randomly sampled a time between 0 and $1$, and masked the sequence by sampling from the forward noising process up to that time. The model was again trained for $400$ epochs without any regularization, at which point the loss function converged. Training the model in this way, wherein the parameters for the non-mask terms are frozen, has the nice property that the noisy model is gauranted to be equivalent to the non-noisy model at $t=1$.
1.11.7 Second-round guided ABE library design
We generated the second-round library by guiding ESM3 and FMIF respectively with the classifier we trained based on the experimental data collected on sequences from the first round. We used the same sampling hyperparameters for each model as we did in the first round. We used exact guidance with guidance strength $\gamma=100$. We designed 944 sequences with each model. All of the generated sequences had a predicted probability of being highly active greater than 0.5. We also included 112 sequences from the first-round library as controls so we could calibrate the log-enrichment measurements between the two rounds. We selected these 112 control sequences with by sampling uniformly across the range of observed log enrichment in the first round.
The 1,888 designed sequences and 112 control sequences were assayed and sequenced with the same protocols as were done in the first round (Supplementary Figure S9). To enable direct comparison between the first- and second-round libraries, we included 112 sequences from the first round as controls in the second-round library (Supplementary Figure S10a). These control sequences were selected to span the full range of observed log enrichment values from the first round, providing calibration points across the entire activity spectrum. This design allowed us to quantify and correct for batch effects between the two experimental rounds. We used these control sequences to learn a linear transformation that maps log enrichment values from the second round to their corresponding values in the first round. Specifically, we fit a linear regression model:
$\log e_{\text{round 1}}=\alpha\cdot\log e_{\text{round 2}}+\beta$ (S36)
@@ 1.11.7 Second-round guided ABE library design
To assess the uncertainty in this calibration procedure, we employed a bootstrap approach using the 112 control sequences. Specifically, we performed 1,000 bootstrap iterations, each time randomly sampling 112 sequences with replacement from our control set and fitting a linear regression model. For each bootstrap sample, we derived a unique set of mapping parameters ($\alpha_{i}$, $\beta_{i}$) and used these to transform the entire second-round dataset. We then calculated the median log enrichment value of the transformed second-round sequences for each bootstrap iteration, generating a distribution of potential outcomes that reflects the inherent variability in our experimental measurements. From this distribution of bootstrap outcomes, we identified the 5th and 95th percentile mappings based on the median log enrichment values. These percentiles represent conservative and optimistic estimates, respectively, of the transformation from round 2 to round 1 log enrichment values (Supplementary Figure S10b).
2 Supplementary Text
2.1 Detailed discussion on related work
Here we provide a more in-depth discussion on the related work to this work. Since ProteinGuide brings technical innovation that are broadly applicable to generative models to the problems in protein engineering, we broadly categorized the related literature into two categories: (i) those that are more related to the technical underpinning of guidance (see subsection “Related work to guidance”), and (ii) those that are more related to the use of pre-trained models and experimental data in protein engineering (see subsection “Related work in protein engineering”). For related work that are more specific to the results showing the equivalence of sampling methods used for discrete flow matching and AO-ARMs, see the Methods section “Simplified and faster sampling for flow matching and diffusion models with masked noise processes”.
2.1.1 Related work to guidance
Several recent work have proposed related but different techniques from this work for steering discrete state-space diffusion models. These approaches can be broadly categorized into on-the-fly approaches (which modify the sampling process at inference time) and training-based approaches (which update the parameters of the pre-trained model).
On-the-fly approaches. Historically, the general problem of sampling from challenging distributions has also been tackled through Markov chain Monte Carlo (MCMC) methods, and some MCMC-based methods have also been proposed for guiding discrete-space generative models. Emami et al. [30] proposed PPDE, which uses the Metropolis–Hastings (MH) algorithm to sample from product-of-experts (PoE) distributions on protein sequence space, where the target is defined via an (unnormalized) sequence-level score combining an “evolutionary density” expert and one or more property experts. The key distinction between that approach to ProteinGuide are that PPDE requires taking gradients through the pre-trained model, which may be prohibitively large; it requires computing (unnormalized) likelihood from the generative model, which is very expensive for most PLMs; and, as it can be quite sensitive to its initialization, potentially having a mixing time that scales exponentially in the sequence length. In contrast, ProteinGuide does not require gradients (or optionally through the predictor, does not require likelihood computations, runs in linear time in sequence length, and does not require a wild-type sequence to initialize sampling. We do note that PPDE can be a useful technique to refine generations from ProteinGuide, particularly when one desires additional local improvements from a good starting point (e.g., a wild-type or a strong candidate) after generating more diverse candidates with ProteinGuide.
Some recent works [67, 99] have also proposed techniques for guiding discrete-space diffusion models using Sequential Monte Carlo (SMC), building upon works first proposed for continuous-space diffusion models [122]. Importantly, ProteinGuide is also complementary to SMC-based approaches as these techniques can be combined with ProteinGuide to improve generation with some additional inference-time costs, such as by way of using multiple particles with proposal distributions defined via TAG. Indeed, through the lens of MCMC, one can also view ProteinGuide as a way to design proposal distributions with minimal rejections by exploiting the structures of the underlying generative process, in a similar spirit to other discrete-space MCMC methods [41]. Lee et al. [65] observed that when applying stronger guidance, sampling can also be improved with SMC, aligning with challenges in low-temperature sampling noted for continuous-space diffusion models [29, 56]. For autoregressive (AR) models, Zhao et al. [131] proposed an SMC-based steering method that learns intermediate twist functions via contrastive learning, similar to learning time-dependent predictors in diffusion guidance.
Vignac et al. [113] propose DiGress, approximate guidance approach in a discrete-time, discrete state-space diffusion (D3PM) [4] framework, which Wang et al. [118] (DPLM) apply to generate protein sequences. In our earlier work [82], we proposed methods for both exact and Taylor-approximate Guidance (TAG), the second of which bears resemblance to DiGress due to both of them using a Taylor series approximation. In our earlier work [82], we discussed the difference between DiGress/DPLM and Discrete Guidance in detail, and performed comprehensive comparisons to DiGress in three experiments across different domains, generally showing Discrete Guidance to be superior. Here we briefly discuss the distinction and refer the readers to Nisonoff et al. [82] (e.g., Appendix section E.2) for more details. DiGress is fundamentally a method for discrete-time, discrete-space diffusion models. It does not specify how to guide either continuous-time, discrete-space diffusion models or discrete-space flow-matching models. ProteinGuide naturally encompasses these model classes, and through our equivalence can be applied to any-order autoregressive models (AO-ARMs). Both theoretically and empirically, the continuous time training objective has been demonstrated to perform better [12, 74, 13, 32]. In fact, one could view DiGress as a special case of our general framework by taking a specific discrete time approximation to the continuous time process. We note that this approximate imposes a uniform time grid and is therefore distinct from our construction of equivalent discrete time processes for continuous time models. Accordingly, their formulation would not directly be applicable to an AO-ARM without additional assumptions. Orthogonally, Yang and Klein [126] proposed FUDGE for guiding autoregressive models, which can be viewed as a special case of our exact guidance method in the case of AR model, with AO-ARMs generalizing AR models.
As noted in Schiff et al. [96], which also presented a discrete-time analog of the continuous-time guidance method in Nisonoff et al. [82], in discrete time, the normalizing constant in Equation S12 is still intractable, necessitating additional assumptions to achieve tractability, whereas Nisonoff et al. [82] showed TAG does not require this assumption. In addition, the continuous-time formulation of discrete diffusion models provides exactly the tractability necessary to perform exact guidance, whereas DiGress is fundamentally approximate. ProteinGuide enables exact guidance both for general noising processes [82] and for masking noise process. In particular, we showed that the model classes discussed—masked diffusion/flow models, masked language models (MLMs), and any-order autoregressive models (AO-ARMs)—are not only equivalent in their training objective, but also the ways they can be sampled (see Methods section “Simplified and faster sampling for flow matching and diffusion models with masked noise processes”). Through this same equivalence, we develop a simpler and faster algorithm for sampling sequences from diffusion and flow-matching models that use masking noise processes, leading to a more computationally efficient guidance algorithm than the one originally proposed in Nisonoff et al. [82], which we called discrete-time exact guidance (DEG). The original method was already shown to at times improve performance over TAG, but we further employed DEG to our in silico experiments to illustrate its effectiveness (as seen in the enzyme and multi-property guidance experiments, where TAG was less effective).
There is another line of work that model discrete data by embedding the data into continuous state-spaces, either on a probabilistic simplex [5, 103] or on a continuous latent space [66, 27]. Some of the approaches proposed to apply guidance to diffusion models on the continuous state-space representations of discrete state-space objects in order to utilize the guidance approaches for continuous state-space diffusion. We note that approaches which perform guidance in continuous spaces lose the discrete structure of the data during generation, which can be important for settings where the discrete structures contain information that are useful for guidance [13]. For more detailed discussion for other inference-time techniques, we refer the readers to a recent review article by Uehara et al. [111].
Training-based approaches. Recent works have also explored updating model parameters to incorporate conditioning rather than modifying the sampling process. Wang et al. [117] proposed to use the Gumbel-Softmax trick to make discrete diffusion trajectories differentiable and enable backpropagation of rewards through entire sampling paths. Concurrently, Rector-Brooks et al. [89] proposed to frame the steering problem as sampling from a Bayesian posterior where the pre-trained model serves as the prior and the reward acts as the likelihood. Both fine-tuning approaches established connections to our earlier work [82] while offering complementary perspectives on solving the steering problem. Direct Preference Optimization (DPO [87]) and related work [120, 18] enable updating of pre-trained models, although up until recently, these methods have been applicable only to models with tractable likelihoods, thereby restricting them largely to autoregressive models. Further research is needed to understand and improve the quality of such approximations.
Reinforcement Learning from Human Feedback (RLHF) offers an alternative approach to model alignment through methods like Proximal Policy Optimization (PPO) [97]. PPO does not technically require the retrained model to have tractable likelihoods, but it requires an explicit reward model and on-policy training, making it often more complicated to train than more direct approaches like DPO [87]. In addition, the standard RLHF pipeline typically begins with fine-tuning, for which the statistical outcome is not well-defined, and for which important information from the already-trained model may be forgotten. For a more detailed review of training-based methods, specifically those anchored in reinforcement learning and diffusion models, we refer the readers to a review article by Uehara et al. [110].
2.1.2 Related work in protein engineering
With the increasing interests of using generative models—both for structures [119, 56] and for sequences [20, 44, 1, 76, 53]—in protein engineering, many of the methods mentioned above have seen adaptations to protein engineering problems, some of which are summarized in a recent review article by Stocco et al. [105]. Below we categorized these applications by the types of methods they employed.
Recent work by Yang et al. [125] benchmarked Discrete Guidance [82] when used with family-specific pre-trained models on a set of protein fitness optimization tasks. They found that Discrete Guidance outperforms several fine-tuning and other “plug-and-play” guidance methods (e.g., methods that perform guidance in continuous latent spaces). In addition, they found that guidance-based approaches require less hyperparameter tuning and have lower computational costs, making them more practically accessible to everyday users. However, these earlier work did not address how to perform guidance for models other than diffusion and flow matching, thereby limiting its utility to the field of protein engineering and beyond. In contrast, by unifying a broad class of discrete-space generative models under a single framework, we are able to directly apply guidance to popular pre-trained models commonly used by practitioners. Accordingly, we do not compare to many of the methods benchmarked in Yang et al. [125] again, but instead focus on demonstrating the efficacy of ProteinGuide on a wider breadth of challenging design scenarios and on a real design campaign.
On-the-fly approaches. In the context of protein sequence generative models, there are prior work that embed protein sequences into continuous state-spaces and thereby directly leveraging guidance techniques developed for continuous-space diffusion models. For example, Gruver et al. [42] propose to perform guidance in the hidden layers of the unconditional diffusion model by jointly training the unconditional denoising model and the classifier on one hidden layer of the denoising model. In another example, Lisanza et al. [71] propose to fine-tune RoseTTAFold with Gaussian diffusion on the one-hot encoded sequence space in order to perform guidance in this continuous state-space.
Another style of approaches that have been very recently developed in the context of large language models involve directly manipulating the intermediate outputs of the model (e.g., activations) [109], and it has only very recently been adapted to protein language models (PLMs) to steer generations toward targeted protein properties [55]. Concretely, these methods compute a “steering direction” in representation space (e.g., from differences in hidden activations associated with an attribute) and then inject this direction into selected layers during the forward pass to bias the model’s next-token (or sequence) distribution, without updating model parameters. Because this line of work is quite recent and protein properties can depend on highly nonlocal and context-dependent constraints, its robustness and generality across objectives, proteins, and model families remains an active area of research.
Training-based approaches. DPO has been applied to a few autoregressive protein generative models. For example, Widatalla et al. [120] applied DPO to ESM-IF1, an autoregressive inverse folding model, to incorporate experimental stability data. In another example, Stocco et al. [104] applied DPO to an autoregressive model conditionally trained on enzyme classes to guide the generation towards various enzyme properties. As for RLHF-based approach, Blalock et al. [8] recently applied PPO to align protein language models (e.g., ESM2) with experimental measurements. Notably, this approach does not directly sample from generative models in the conventional sense; instead, it employs a position-wise scoring strategy where models evaluate mutation probabilities at individual positions sequentially, followed by an iterative selection process that combines high-scoring mutations.
Zero-shot scoring-based approaches. A complementary line of work uses pre-trained protein language models primarily as scoring functions to guide iterative mutation and selection, rather than sampling from an explicit conditional generative model. For example, EVOLVEpro [59] and related methods use language-model-derived scores within iterative optimization or active-learning-style loops to propose and evaluate candidate mutations. Similarly, Hie et al. [45] demonstrate efficient evolution of human antibodies using general protein language models, emphasizing iterative selection guided by model scores rather than direct conditional sampling from a guided generative process. These approaches are highly practical for local improvement starting from strong initial sequences, but are conceptually distinct from plug-and-play conditional sampling methods that directly reweigh the generative trajectory.
Supervised fitness optimization. A large body of work in machine-learning-assisted directed evolution trains supervised predictors $p(y\mid x)$ on experimentally measured variants and then uses these predictors to select new variants to test, often in active learning or Bayesian optimization style loops. Examples include ML-assisted directed evolution with combinatorial libraries [123] and subsequent work emphasizing the importance of training set design and “holey” fitness landscapes [121], and has explored in more data-efficient regime [7]. These methods can be viewed as attempting to invert a learned predictor to propose improved sequences, but unlike generative-model-based approaches they do not provide an explicit conditional sequence distribution and typically rely on local search or proposal mechanisms.
2.2 Detailed derivations and proofs
Below we included the detailed derivations and proofs for the statements and lemmas in the Methods section “Statement of exact MFM sampling using AO-ARMs”.
2.2.1 Proofs of exact MFM sampling using AO-ARMs
Proof of Proposition 1.
Since CTMCs are, by definition memoryless (depend only on the previous time and state), the path probability density decomposes into a product of densities for each transition conditioned on the last. This can be written as
$\mathbb{P}_{\mathrm{MFM}}\left(\bm{\tau},\bm{X}\right)$
$=$ $\mathbb{P}_{\mathrm{MFM}}(\tau_{0},x_{\tau_{0}})\times\prod_{i=1}^{D}\mathbb{P}_{\mathrm{MFM}}(\tau_{i},x_{\tau_{i}}|\tau_{i-1},x_{\tau_{i-1}})\times\mathbb{P}_{\mathrm{MFM}}(\tau_{D+1}>1|\tau_{D},x_{\tau_{D}}).$
The MFM always starts at time $\tau_{0}=0$ with the fully masked sequence $x_{0}=\{M\}^{D}$ so $\mathbb{P}_{\mathrm{MFM}}(\tau_{0},x_{\tau_{0}})=1$. There are also exactly $D$ transitions because we assume 0 stochasticity and no corrector sampling. This means $\mathbb{P}_{\mathrm{MFM}}(\tau_{D+1}>1|\tau_{D},x_{\tau_{D}})=1$ as well. The above now simplifies to
$=$ $\prod_{i=1}^{D}\mathbb{P}_{\mathrm{MFM}}(\tau_{i},x_{\tau_{i}}|\tau_{i-1},x_{\tau_{i-1}})$
$=$ $\prod_{i=1}^{D}\mathbb{P}_{\mathrm{MFM}}(\tau_{i}|\tau_{i-1},x_{\tau_{i-1}})\mathbb{P}_{\mathrm{MFM}}(x_{\tau_{i}}|\tau_{i},x_{\tau_{i-1}}).$
We can apply Lemma 1 to remove the state dependence from the conditional distribution over jump times, $\mathbb{P}_{\mathrm{MFM}}(\tau_{i}|\tau_{i-1},x_{\tau_{i-1}})$. Similarly, we can apply Lemma 2 to remove the time dependence from the conditional distribution over jump states, $\mathbb{P}_{\mathrm{MFM}}(x_{\tau_{i}}|\tau_{i},x_{\tau_{i-1}})$. This is the key step of the proof, allowing us to separate the time- and state-dependent distributions into separate terms as follows
$=$ $\prod_{i=1}^{D}\mathbb{P}_{\mathrm{MFM}}(\tau_{i}|\tau_{i-1})\mathbb{P}_{\mathrm{MFM}}(x_{\tau_{i}}|x_{\tau_{i-1}})$
$=$ $\left(\prod_{i=1}^{D}\mathbb{P}_{\mathrm{MFM}}(\tau_{i}|\tau_{i-1})\right)\left(\prod_{i=1}^{D}\mathbb{P}_{\mathrm{MFM}}(x_{\tau_{i}}|x_{\tau_{i-1}})\right).$
In Lemmas 1 and 2 we also calculate closed-form expressions for $\mathbb{P}_{\mathrm{MFM}}(\tau_{i}|\tau_{i-1})$ and $\mathbb{P}_{\mathrm{MFM}}(x_{\tau_{i}}|x_{\tau_{i-1}})$, which can be substituted in and simplified as follows
$=$ $\prod_{i=1}^{D}\dot{\kappa}_{\tau_{i}}\times\prod_{i=1}^{D}p_{1|\tau_{i}}(X_{1}=x_{\tau_{i}}|X_{\tau_{i}}^{-}=x_{\tau_{i-1}}).$
Now, we will rewrite these products to directly match the proposition statement. First, the interpolation schedule $\kappa_{t}$ is the CDF of the jump time under the assumption of minimal transitions. This means the PDF of the jump time for a given position is $\mathbb{P}_{\mathrm{MFM}}(\tau=t)=\frac{d}{dt}\mathbb{P}_{\mathrm{MFM}}(\tau<t)=\dot{\kappa}_{t}$. Generally, we choose $\kappa_{t}=t$ so the PDF of the jump times is uniform $\mathbb{P}_{\mathrm{MFM}}(t=\tau)=\dot{\kappa}_{t}=1=U_{[0,1]}(t=\tau)$:
$=$ $\prod_{i=1}^{D}U[0,1](\tau_{i})\times\prod_{i=1}^{D}p_{1|\tau_{i}}(X_{1}=x_{\tau_{i}}|X_{\tau_{i}}^{-}=x_{\tau_{i-1}}).$
Finally, we recognize that the AO-ARM constructed for the sampling algorithm uses the MFM denoiser as its conditional distribution. This means the second product above is just the conditional probability of the sample under the AO-ARM   and the expression simplifies to
$=$ $\prod_{i=1}^{D}U[0,1](\tau_{i})\times\prod_{i=1}^{D}p(X^{\sigma(i)}=x_{\tau_{i}}|X^{\sigma(<i)}=x_{\tau_{i-1}})$
$=$ $U^{D}[0,1](\bm{T}=\sigma^{-1}(\bm{\tau}))\times\mathbb{P}_{\mathrm{OA}}(X=X_{\tau_{D}}|\sigma)$
as claimed. ∎
Proof of Corollary 1.
This follows directly from Proposition 1. Originally, we found
$\mathbb{P}_{\mathrm{MFM}}(\tau,\bm{X})$ $=\mathbb{P}_{\mathrm{OA}}(X_{\tau_{D}}|\sigma)U^{D}[0,1](\bm{T}=\sigma^{-1}(\bm{\tau}))$
To be able to sample both the times and permutations independently, upfront, our sampler must find the permutation $\pi$ that maps the sampled transition times $\bm{T}$ to sampled positions. Then applying $\sigma$ to $\pi(T)$ sorts the times to be $\bm{\tau}$. In practice, we would just assign the sampled times to positions in the order specified by $\sigma$. For the purposes of this proof, however, $\pi$ lets us rewrite this expression as:
$=U[S_{D}](\sigma)\mathbb{P}_{\mathrm{OA}}(X_{\tau_{D}}|\sigma)U^{D}[0,1](\bm{T}=\pi^{-1}(\sigma^{-1}(\bm{\tau})))$
Finally, we apply the definition of OA-AR sampling
$=\mathbb{P}_{\mathrm{OA}}(\bm{X})U^{D}[0,1](\bm{T})$
2.2.2 Statements and proofs of lemmas
Lemma 1.
The conditional probability of the next jump time for an MFM depends only on the previous jump time and can be expressed in closed form as:
$\mathbb{P}_{\mathrm{MFM}}\left(\tau_{i}|\tau_{i-1},x_{\tau_{i-1}}\right)$ $=\mathbb{P}_{\mathrm{MFM}}\left(\tau_{i}|\tau_{i-1}\right)$
$=(D-(i-1))\frac{\dot{\kappa}_{\tau_{i}}}{1-\kappa_{\tau_{i-1}}}\left(\frac{1-\kappa_{\tau_{i}}}{1-\kappa_{\tau_{i-1}}}\right)^{D-i}.$ (S37)